### v1

@@ Abstract
Figure 4 demonstrates that the compressed eBOSS Lyman-alpha forest likelihood is accurate for the cosmological models that we consider given the precision of eBOSS data. We take the best-fit linear matter power spectra for each model and rescale them such that they have the same values of $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}}$ as the best-fit $\Lambda$CDM model. Once the information in these two parameters is accounted, there is negligible difference between the model power spectra compared to the $\sim 6\%$ precision of eBOSS data (conservatively estimated from the constrained uncertainty on $\Delta_{\mathrm{lin}}^{2}$). The range of wavenumbers to which eBOSS is sensitive is estimated from the velocity wavenumbers measured in the flux power spectrum, although the flux power spectrum is in fact sensitive to all wavenumbers since it is an integral over modes transverse to the line of sight. The validity of the compression is nonetheless confirmed by explicitly demonstrating that the posterior distribution is identical whether inferred using the full flux power spectrum or the compressed likelihood [29]. Given the anticipated precision of upcoming DESI data, it may be necessary to add a parameter to the compression capturing curvature in the power spectrum, in particular to account for ULA models. This extra information may also distinguish further between cosmological models.
@@
$\log[m_{\mathrm{WDM}}\,({\rm eV})]$ Unconstrained (0.21) 1.35($1.35(1.95)^{+0.60}_{-0.30}$
$\frac{T_{\mathrm{WDM}}}{T_{\mathrm{CMB}}}$ $<0.168(0.001)$ 0.155($0.155(0.194)^{+0.018}_{-0.052}$
$f_{\rm WDM}$ $<0.048(0.005)$ 0.0250($0.0250(0.0219)^{+0.0021}_{-0.0079}$
$h$ $0.6744(0.6745)\pm 0.0041$ $0.6745(0.6748)\pm 0.0040$
$\omega_{\mathrm{CDM}}$ 0.1187($0.1187(0.1198)^{+0.0019}_{-0.0010}$ 0.1175($0.1175(0.1178)^{+0.0014}_{-0.0011}$
$10^{2}\omega_{\mathrm{b}}$ $2.236(0.022)\pm 0.013$ $2.236(0.022)\pm 0.013$
$10^{9}A_{\mathrm{s}}$ $2.106(2.104)\pm 0.030$ $2.100(2.104)\pm 0.030$
$n_{\mathrm{s}}$ $0.9646(0.9657)\pm 0.0037$ $0.9642(0.9655)\pm 0.0036$
$\tau$ 0.0561($0.0561(0.0555)^{+0.0068}_{-0.0077}$ $0.0548(0.0554)\pm 0.0070$
$\Delta_{\mathrm{lin}}^{2}$ 0.3451($0.3451(0.3491)^{+0.0095}_{-0.0055}$ 0.331($0.331(0.322)^{+0.014}_{-0.0085}$
$n_{\mathrm{lin}}$ 2.315($-2.315(-2.304)^{+0.015}_{-0.0009}$ $-2.3365(-2.3382)\pm 0.0058$
$S_{8}$ $0.830(0.830)\pm 0.010$ $0.824(0.821)\pm 0.011$
$\Omega_{\mathrm{m}}$ $0.3142(0.3139)\pm 0.0056$ $0.3140(0.3136)\pm 0.0055$
@@
Full sets of marginalized posteriors – Figures 5 to 9 show full sets of 1D and 2D marginalized posteriors of cosmological parameters given the different cosmological models and data combinations. Fig. 5 shows the case for the $\Lambda$CDM model. The two-parameter compressed eBOSS Lyman-alpha forest-only likelihood is, unsurprisingly, unable to constrain the six-parameter baseline cosmology. Even after adding information through a Gaussian prior on $H_{0}$ inspired by Planck CMB results ($H_{0}\sim\mathcal{N}(67.3,1)\,\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}$) as done by the eBOSS Collaboration, the other standard $\Lambda$CDM parameters remain poorly constrained. This is different from the results presented by the eBOSS Collaboration who find that, after additionally fixing the baryon energy density, they simultaneously constrain $\sigma_{8}$, $n_{\mathrm{s}}$ and the total matter energy density $\Omega_{\mathrm{m}}$ using a full flux power spectrum likelihood [26].
@@
However, this contradicts the results of Refs. [27, 28, 29, 30] which demonstrate that full flux power spectrum and compressed likelihoods return identical posteriors on $\Lambda$CDM parameters – and the re-analysis of eBOSS flux power spectra [51] that constrains two cosmological parameters (the amplitude and tilt of the primordial power spectrum defined at $k=0.78h\,\mathrm{Mpc}^{-1}$) after using a Gaussian $H_{0}$ prior, a uniform prior on $\Omega_{\mathrm{m}}h^{2}$ restricting it to values compatible with Planck and fixing $\omega_{\mathrm{b}}$. In testing, we find that the eBOSS-only $\Lambda$CDM constraints are highly sensitive to the width of uniform priors on cosmological parameters. Indeed, we recover the same degeneracy as the eBOSS Collaboration between $n_{\mathrm{s}}$ and $\Omega_{\mathrm{m}}$ and the same mild discrepancy with Planck contours in this plane, except that the constraint along the degeneracy direction is prior-dependent. The full flux power spectrum likelihood used by the eBOSS Collaboration is not publicly available and so we defer to future work a detailed comparison of eBOSS-only constraints, in particular accounting for the effect of prior distributions.
@@
Our measurement of tension in the $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}}$ parameters is in exact agreement with results of the eBOSS Collaboration as we use the compressed likelihood as calculated by them. Further, we reproduce the combined Planck, BAO and eBOSS constraints in the $\Lambda$CDM and varying neutrino mass models (Fig. 6) where we can compare. This result arises since, once data are sufficiently constraining on the full six- or seven-parameter models, the posterior loses sensitivity to the choice of prior. Therefore, our primary conclusions regarding the tension in the $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}}$ parameters and combined CMB, BAO, SNe and Lyman-alpha forest constraints on extended models is unaffected by the choice of prior or the use of a full or compressed likelihood.

### v3

@@ Abstract
Figure 4 demonstrates that the compressed eBOSS Lyman-alpha forest likelihood is accurate for the cosmological models that we consider given the precision of eBOSS data. We take the best-fit linear matter power spectra for each model and rescale them such that they have the same values of $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}}$ as the best-fit $\Lambda$CDM model. Once the information in these two parameters is accounted, there is negligible difference between the model power spectra compared to the $\sim 6\%$ precision of eBOSS data (conservatively estimated from the constrained uncertainty on $\Delta_{\mathrm{lin}}^{2}$). The range of wavenumbers to which eBOSS is sensitive is estimated from the velocity wavenumbers measured in the flux power spectrum, although the flux power spectrum is theoretically sensitive to all wavenumbers since it is an integral over modes transverse to the line of sight. The range is calculated by Ref. [22] which demonstrates that the sensitivity of the eBOSS 1D flux power spectrum to the linear matter power spectrum is negligible for smaller or larger wavenumbers given the poor observational sampling of modes transverse to the line of sight. To be additionally conservative, we confirm that our extended models remain sufficiently close to a power law over an even longer range of wavenumbers $0.1\,h\,\mathrm{Mpc}^{-1}<k<8\,h\,\mathrm{Mpc}^{-1}$. The validity of the compression is further confirmed by explicitly demonstrating that the posterior distribution is identical whether inferred using the full flux power spectrum or the compressed likelihood [30].
The compression is invalid for models with sharp features in the matter power spectrum, e. g., ULA or WDM models where the new particle forms the entirety of the dark matter. We deliberately restrict ourselves to mixed dark matter models, with a prior set to preclude these features forming. Given the anticipated precision of upcoming DESI data, it may be necessary to add a parameter to the compression capturing curvature in the power spectrum, in particular to account for mixed ULA models. This extra information may also distinguish further between cosmological models.
@@
$\log[m_{\mathrm{WDM}}\,({\rm eV})]$ Unconstrained (0.21) $1.01(0.95)^{+0.30}_{-0.44}$
$\frac{T_{\mathrm{WDM}}}{T_{\mathrm{CMB}}}$ $<0.261(0.001)$ 0.194($0.194(0.194)^{+0.043}_{-0.074}$
$f_{\rm WDM}$ $<0.049(0.005)$ 0.0219($0.0219(0.0219)^{+0.0030}_{-0.0042}$
$h$ $0.6745(0.6745)\pm 0.0041$ $0.6741(0.6748)\pm 0.0040$
$\omega_{\mathrm{CDM}}$ 0.1189($0.1189(0.1198)^{+0.0018}_{-0.0009}$ $0.1180(0.1178)\pm 0.0011$
$10^{2}\omega_{\mathrm{b}}$ $2.237(2.237)\pm 0.014$ $2.237(2.237)\pm 0.013$
$10^{9}A_{\mathrm{s}}$ $2.105(2.104)\pm 0.030$ $2.105(2.104)\pm 0.029$
$n_{\mathrm{s}}$ $0.9648(0.9657)\pm 0.0037$ $0.9640(0.9655)\pm 0.0036$
$\tau$ $0.0558(0.0555)\pm 0.0072$ $0.0558(0.0554)\pm 0.0068$
$\Delta_{\mathrm{lin}}^{2}$ 0.341($0.341(0.349)^{+0.014}_{-0.005}$ 0.324($0.324(0.322)^{+0.013}_{-0.010}$
$n_{\mathrm{lin}}$ 2.316($-2.316(-2.304)^{+0.015}_{-0.0009}$ $-2.3367(-2.3382)\pm 0.0058$
$S_{8}$ $0.827(0.830)\pm 0.011$ $0.822(0.821)\pm 0.012$
$\Omega_{\mathrm{m}}$ $0.3141(0.3139)\pm 0.0055$ $0.3148(0.3136)\pm 0.0055$
@@
Other likelihoods and analysis details – For CMB data, we use the Plik low-multipole $\ell$ TT and EE, $-\ell$ TTTEEE and conservative lensing $\phi\phi$ likelihoods from the Planck 2018 data release [1, 3]. We use BAO measurements from the CMASS and LOWZ galaxy samples of BOSS Data Release 12 at $z=0.38$, $0.51$ and $0.61$ [33], the 6dF Galaxy Survey (6dFGS) at $z=0.106$ [34] and SDSS Data Release 7 Main Galaxy Sample (SDSS DR7 MGS) at $z=0.15$ [35]; and the Pantheon+ catalog of uncalibrated SNe luminosity distances [7]. This compendium of CMB, BAO and SNe data constitutes a larger-scale complement to the smaller scales probed by the eBOSS flux power spectrum. We vary nuisance foreground parameters in the Planck likelihood and SNe calibrations in the Pantheon+ likelihood.
We use a uniform prior distribution on baseline $\Lambda$CDM parameters: the Hubble constant $H_{0}$ [^6] For technical reasons, we sometimes instead use the angular size of the sound horizon $\theta_{\mathrm{s}}$., the physical baryon $\omega_{\mathrm{b}}$ and CDM $\omega_{\rm CDM}$ energy densities, the amplitude $A_{\mathrm{s}}$ and tilt $n_{\mathrm{s}}$ of the primordial power spectrum and the reionization optical depth $\tau$. For the extended cosmological parameters, we set uniform prior ranges $\mathrm{log}[m_{\rm ULA}\,(\mathrm{eV})]\in[-27,-23],f_{\rm ULA}\in[0,0.05],\mathrm{log}[m_{\rm WDM}\,(\mathrm{eV})]\in[0,3],\frac{T_{\rm WDM}}{T_{\rm CMB}}\in[0,1]$, where $T_{\mathrm{CMB}}$ is the CMB temperature, with the additional constraint $f_{\rm WDM}<0.05$. Unless specified otherwise, we model free-streaming neutrinos as two massless and one massive neutrino with the sum of neutrino masses $\Sigma m_{\nu}=0.06$ eV.
We use the cosmological inference package MontePython-v3777https://github.com/brinckmann/montepython_public. [70, 71] interfaced with a modified version [^8] https://github.com/PoulinV/AxiCLASS. of the Einstein-Boltzmann solver CLASS [^9] https://lesgourg.github.io/class_public/class.html. [72, 73] in order to perform a Markov chain Monte Carlo (MCMC) sampling of the posterior distribution using a Metropolis Hasting algorithm. We model ULAs following the effective fluid approach from the modified Einstein-Bolztmann solver AxionCAMB [74] and implemented in AxiCLASS [^10] An improved fluid model was derived in Ref. [75] but the difference is negligible given the precision of our data. [36, 37]. We consider MCMC chains converged when the Gelman-Rubin statistic $|R-1|\lesssim 0.05$ [76]. To analyze MCMC chains and visualize posteriors, we use the GetDist package [77]. We find minimum $\chi^{2}$ values using the simulated annealing method described in the appendix of Ref. [78].
@@
Full sets of marginalized posteriors – Figures 5 to 9 show full sets of 1D and 2D marginalized posteriors of cosmological parameters given the different cosmological models and data combinations. Fig. 5 shows the case for the $\Lambda$CDM model. The two-parameter compressed eBOSS Lyman-alpha forest-only likelihood is, unsurprisingly, unable to constrain the six-parameter baseline cosmology. Even after adding information through a Gaussian prior on $H_{0}$ inspired by Planck CMB results ($H_{0}\sim\mathcal{N}(67.3,1)\,\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1}$) as done by the eBOSS Collaboration, the other standard $\Lambda$CDM parameters remain poorly constrained. This is different from the results presented in Ref. [27]: after additionally fixing the baryon energy density, they simultaneously constrain $\sigma_{8}$, $n_{\mathrm{s}}$ and the total matter energy density $\Omega_{\mathrm{m}}$ using a full flux power spectrum likelihood.
@@
However, this contradicts the results of Refs. [28, 29, 30, 31] which demonstrate that full flux power spectrum and compressed likelihoods return identical posteriors on $\Lambda$CDM parameters – and the re-analysis of eBOSS flux power spectra [46] that constrains two cosmological parameters (the amplitude and tilt of the primordial power spectrum defined at $k=0.78h\,\mathrm{Mpc}^{-1}$) after using a uniform prior on $\Omega_{\mathrm{m}}h^{2}$ restricting it to values compatible with Planck and fixing $\omega_{\mathrm{b}}$ [^11] Ref. [46] constrains $H_{0}$ as an artifact of converting from velocity to spatial units in a finite simulation box, i. e., there is no physical information on $H_{0}$ in the data.. In testing, we find that the eBOSS-only $\Lambda$CDM constraints are highly sensitive to the width of uniform priors on cosmological parameters. Indeed, we recover the same degeneracy as the eBOSS Collaboration between $n_{\mathrm{s}}$ and $\Omega_{\mathrm{m}}$ and the same mild discrepancy with Planck contours in this plane, except that the constraint along the degeneracy direction is prior-dependent. The full flux power spectrum likelihood used by the eBOSS Collaboration is not publicly available and so we defer to future work a detailed comparison of eBOSS-only constraints, in particular accounting for the effect of prior distributions.
@@
Our measurement of tension in the $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}}$ parameters is in exact agreement with results of the eBOSS Collaboration as we use the compressed likelihood as calculated by them. Further, we reproduce the combined Planck, BAO and eBOSS constraints in the $\Lambda$CDM and varying neutrino mass models (Fig. 6) of Ref. [27]. This result arises since, once data are sufficiently constraining on the full six- or seven-parameter models, the posterior loses sensitivity to the choice of prior. Therefore, our primary conclusions regarding the tension in the $\Delta_{\mathrm{lin}}^{2},n_{\mathrm{lin}}$ parameters and combined CMB, BAO, SNe and Lyman-alpha forest constraints on extended models is unaffected by the choice of prior or the use of a full or compressed likelihood.